# Allocation Multiplicity: Evaluating the Promises of the Rashomon Set

### v1

@@ 4.2. From the Perspective of the Rashomon Set
Prior works suggest that the Rashomon set promises benefits such as less discriminatory algorithms, reduced outcome homogenization, and fairer decisions through ensembles or reconciliation. However, we contend that allocations derived from the Rashomon set (e.g., via a top-$k$ mapping) may fall short of fully delivering on these promises. To the extent that the Rashomon set helps us deliver these benefits, it is because it helps us to recover more allocations from the broader space of equal-utility allocations. In this section, we argue that considering equal-utility allocations directly can better achieve these aims.
@@
For a given fairness metric, there exists a least discriminatory allocation within the space of all equal-utility allocations. For example, decision-makers may accidentally under-select individuals who are racial minorities because they are perceived as riskier by predictive models. However, excluding these individuals is not necessary: provided there are at least $k^{\prime}$ qualified individuals within the minority population, an equal-utility allocation exists that consists entirely of minority individuals. But since the all, majority, or even proportional minority allocations are difficult for predictive models to find, this may cause all “less discriminatory algorithms” (LDAs) in the Rashomon set to result in allocations that deviate significantly from the “least discriminatory” equal-utility allocation, as seen in Figure 3(a). Moreover, generalization error further contributes to this issue: prior research has shown that LDAs derived from training or validation data can become more discriminatory when deployed in unseen settings (Black et al., 2024a; Laufer et al., 2024).
@@
In the space of all equal-utility allocations, every individual has a non-zero probability of being selected when allocations do not achieve perfect utility. This means that the systemic rejection rate could theoretically be reduced to zero without sacrificing utility, provided there are enough decision-makers. However, as illustrated in Figure 3(b), allocations derived from the Rashomon set may still exclude some qualified individuals entirely, leaving them with zero chance of selection. This underscores that predictive multiplicity, as captured by the Rashomon set, does not fully deliver on its promise of reducing homogenization if this is a primary concern. Viewing the problem through the lens of allocation multiplicity makes this limitation clear and highlights how considering the space of equal-utility allocations can provide better insight into and potential pathways for reducing outcome homogenization.
@@
Ensembling or reconciling predictions across models in the Rashomon set is often proposed as a way to achieve fairer decisions. However, these methods are limited when viewed through the lens of allocation multiplicity. While they can or smooth predictions across models, they may fail to capture the full range of equal-utility allocations that the Rashomon set implicitly supports. Ensembling or reconciling predictions necessarily is based on the starting set of Rashomon predictions, but the Rashomon set is not an even representation of the set of equal-utility allocations. A default decision-making approach—such as using a top-$k$ mapping may not account for the diversity of allocations or represent the average characteristics of individuals chosen across the space of equal-utility allocations. This limitation underscores the need to consider allocation multiplicity directly, as it provides a more comprehensive view of fairness by encompassing all plausible and representative outcomes within the space of equal-utility allocations.
@@ 5.1. Data and Methods
Different Bootstrapped Shuffled Weight
Feature Subsets Training Data Data Order Perturbation
$\frac{n^{\prime}}{n}$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$
Number of Rashomon Models 0.32 1674 1674 997 997 959 959 999 999
0.55 1386 1386 1000 1000 835 835 1000 1000
Validation Loss 0.32 0.279 0.279 0.284 0.284 0.284 0.284 0.290 0.290
0.55 0.327 0.327 0.328 0.328 0.330 0.330 0.335 0.335
Number of Recovered Allocations 0.32 502 373 997 997 75 62 54 68
0.55 734 460 1000 1000 184 74 26 31
Allocation Utility ($k^{\prime}/k$) 0.32 0.873 0.602 0.877 0.604 0.877 0.605 0.877 0.604
0.55 0.950 0.902 0.953 0.904 0.952 0.905 0.952 0.905
@@
Using this data, we run various simulations of a scenario in which a hospital is allocating enrollment in a healthcare program among its patients. In order to determine who is qualified for the program (Obermeyer et al., 2019; Watson-Daniels et al., 2023a), the hospital could predict which patients will have the highest future healthcare costs or the most chronic illnesses. However, Obermeyer et al. (2019) demonstrate that using predicted healthcare costs to select patients can result in racial bias. We therefore consider an individual to be qualified if they will have $q$ or more chronic illnesses in the upcoming year. To test different qualification rates, we vary $q$ to represent having at least 1, 2, or 3 chronic illnesses, which corresponds to 55%, 32%, and 19% of the population, respectively. Not all qualified patients can be selected because, as in the real-world [^8] In the real-world hospital studied by Obermeyer et al. (2019), patients with risk scores in the top 3% were automatically enrolled in the program, while those in the top 45% were flagged for their doctors, who could enroll them based on program availability., healthcare programs are resource-constrained (Obermeyer et al., 2019; Watson-Daniels et al., 2023a). To simulate this, we consider different selection rates of 10%, 25%, and 50%. For each combination of qualification and selection rates, we run 250 simulations [^9] We use 10 different partitions of a 60-20-20 train-validation-test split. For each partition, we run 25 simulations where a different set of $n=1000$ patients are drawn from the test set to be candidates for our hypothetical allocation. with different test sets of $n=1000$ patients as allocation candidates.
@@
(a) The “least discriminatory” equal-utility allocation is not found based on the threshold test, which evaluates if the ratio of chronic illnesses between selected black and white patients is greater than 1.
(b) Outcome homogenization persists in recovered allocations from the Rashomon set: between 15% to 30% of qualified individuals are systemically rejected.
(c) The age of selected patients is skewed older in the average allocation from the Rashomon set, indicating how ensembling predictions may lead to unrepresentative allocations.
@@
$k/n$ $n^{\prime}/n$ All Equal-Utility Different Bootstrapped Shuffled Weight

### v3

@@ 4.2. From the Perspective of the Rashomon Set
Prior works suggest that the Rashomon set promises benefits such as less discriminatory algorithms, reduced outcome homogenization, and fairer decisions through ensembles or reconciliation. In this section, we argue that the space of equal-utility allocations is what actually promises these benefits. The extent to which the Rashomon set fulfills these promises depends on how effectively it recovers diverse equal-utility allocations.
@@
For a given fairness metric, there exists a least discriminatory allocation within the space of all equal-utility allocations. For example, decision-makers may accidentally under-select individuals who are racial minorities because they are perceived as riskier by predictive models. However, excluding these individuals is not necessary: provided there are at least $k^{\prime}$ qualified individuals within the minority population, an equal-utility allocation exists that consists entirely of minority individuals. But since most of these allocations are difficult for predictive models to find, this may cause all “less discriminatory algorithms” (LDAs) in the Rashomon set to result in allocations that deviate significantly from the “least discriminatory” equal-utility allocation, as Figure 3(a) illustrates. Moreover, train-test data differences may further contribute to this issue: prior research has shown that LDAs derived from training or validation data can become more discriminatory when deployed in unseen deployment settings (Black et al., 2024a; Laufer et al., 2024).
@@
In the space of all equal-utility allocations, every individual has a non-zero probability of being selected when allocations do not achieve perfect utility. This means that the systemic rejection rate could theoretically be reduced to zero without sacrificing utility, provided there are enough decision-makers. However, as illustrated in Figure 3(b), allocations derived from the Rashomon set may exclude some qualified individuals entirely, leaving them with zero chance of selection. This underscores that predictive multiplicity, as captured by the Rashomon set, does not fully deliver on its promise of reducing homogenization. Viewing the problem through the lens of allocation multiplicity makes this limitation clear and highlights how considering the space of equal-utility allocations can provide better insights into pathways for reducing outcome homogenization.
@@
Ensembling or reconciling predictions across models in the Rashomon set is often proposed as a way to achieve fairer decisions. However, these methods are limited when viewed through the lens of allocation multiplicity. While they resolve inconsistent predictions across models, the aggregated predictions may not capture the diversity of equal-utility allocations that the Rashomon set implicitly supports. Ensembling or reconciling predictions is necessarily based on the starting set of Rashomon predictions, but allocations yielded by the Rashomon set (via a top-$k$ mapping) may not form an even representation of the set of equal-utility allocations. For example, the average characteristics of individuals chosen across the space of equal-utility allocations may differ greatly from the average characteristics of individuals chosen in allocations recovered by the Rashomon set. This limitation underscores the need to consider the space of equal-utility allocations directly, at it provides a more comprehensive view of fairness by encompassing all diverse outcomes.
@@ 5.1. Data and Methods
Different Bootstrapped Shuffled Weight Decision-Boundary
Feature Subsets Training Data Data Order Perturbation Randomization
$\frac{n^{\prime}}{n}$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$ $\frac{k}{n}=0.25$ $\frac{k}{n}=0.50$
Number of Rashomon Models 0.32 1674 1674 997 997 959 959 999 999 1 1
0.55 1386 1386 1000 1000 835 835 1000 1000 1 1
Validation Loss 0.32 0.279 0.279 0.284 0.284 0.284 0.284 0.290 0.290 0.281 0.281
0.55 0.327 0.327 0.328 0.328 0.330 0.330 0.335 0.335 0.325 0.325
Number of Recovered Allocations 0.32 502 373 997 997 75 62 54 68 1000 1000
0.55 734 460 1000 1000 184 74 26 31 1000 1000
Allocation Utility ($k^{\prime}/k$) 0.32 0.873 0.602 0.877 0.604 0.877 0.605 0.877 0.604 0.869 0.599
0.55 0.950 0.902 0.953 0.904 0.952 0.905 0.952 0.905 0.952 0.885
@@
Using this data, we run various simulations [^9] The code for our experiments is available at: https://github.com/mqwangwa/allocation_multiplicity. of a scenario in which a hospital is allocating enrollment in a healthcare program among its patients. In order to determine who is qualified for the program (Obermeyer et al., 2019; Watson-Daniels et al., 2023a), the hospital could predict which patients will have the highest future healthcare costs or the most chronic illnesses. However, Obermeyer et al. (2019) demonstrate that using predicted healthcare costs to select patients can result in racial bias. We therefore consider an individual to be qualified if they will have $q$ or more chronic illnesses in the upcoming year. To test different qualification rates, we vary $q$ to represent having at least 1, 2, or 3 chronic illnesses, which corresponds to 55%, 32%, and 19% of the population, respectively. Not all qualified patients can be selected because, as in the real-world, [^10] In the real-world hospital studied by Obermeyer et al. (2019), patients with risk scores in the top 3% were automatically enrolled in the program, while those in the top 45% were flagged for their doctors, who could enroll them based on program availability. healthcare programs are resource-constrained (Obermeyer et al., 2019; Watson-Daniels et al., 2023a). To simulate this, we consider different selection rates of 10%, 25%, and 50%. For each combination of qualification and selection rates, we run 250 simulations [^11] We use 10 different partitions of a 60-20-20 train-validation-test split. For each partition, we run 25 simulations where a different set of $n=1000$ patients are drawn from the test set to be candidates for our hypothetical allocation. Results with standard deviation over these 250 iterations are reported in the Appendix. with different test sets of $n=1000$ patients as allocation candidates.
@@
For comparison, we also evaluate stochastically generated allocations using a single model’s predictions and decision-boundary randomization. We follow the procedure in Jain et al. (2024a) to generate allocations using a partial weighted lottery [^13] In the Appendix, we try other parameters for decision-boundary randomization as well as another method to randomize allocations (see Table 8). Generally, randomized methods achieve better results in fulfilling the promises of the Rashomon set while still attaining comparable utility.: the first $0.75k$ resources are deterministically given to patients with the highest predictions and then the remaining $0.25k$ resources are randomized over patients with the next $0.50k$-highest predictions. While these allocations are not recovered from a Rashomon set of models, they represent another way to find (close to) equal-utility allocations. Notably, the mean and variance of deployment utility are similar for recovered allocations from the Rashomon set and those from decision-boundary randomization (Appendix Table 8).
(a) The “least discriminatory” allocation is not found, according to the threshold test and reference equal-utility allocation that selects patients in descending order by number of chronic illnesses.
(b) Outcome homogenization persists in recovered allocations from the Rashomon set: between 10% to 30% of qualified individuals are rejected by all Rashomon models.
(c) The age of selected patients is skewed older on average in recovered allocation from the Rashomon set, indicating how ensembling predictions may lead to unrepresentative allocations.
@@
$k/n$ $n^{\prime}/n$ All Equal-Utility Different Bootstrapped Shuffled Weight Decision-Boundary
Allocations Feature Subsets Training Data Data Order Perturbation Randomization
Least Discriminatory Allocation $\dagger$ 0.25 0.32 0.86 1.12 1.12 1.19 1.18 1.11
0.25 0.55 0.92 1.10 1.12 1.18 1.20 1.14
0.50 0.32 1.21 1.26 1.25 1.31 1.28 1.23
0.50 0.55 0.90 1.25 1.24 1.30 1.29 1.23
Outcome Homogenization $\ddagger$ 0.25 0.32 0.80 0.95 0.97 0.99 0.99 0.94
0.25 0.55 0.71 0.92 0.95 0.98 0.99 0.94
0.50 0.32 0.69 0.95 0.96 0.99 0.98 0.89
0.50 0.55 0.76 0.94 0.96 0.99 0.99 0.89
Age Homogenization $\star$ 0.25 0.32 2.49 2.35 2.33 2.36 2.36 2.38
0.25 0.55 2.56 2.20 2.21 2.31 2.33 2.23
0.50 0.32 2.52 2.51 2.48 2.47 2.47 2.52
0.50 0.55 2.55 2.49 2.46 2.45 2.45 2.50
$\dagger$ The least discriminatory allocation is measured by the lowest found ratio of active chronic illnesses between selected Black and White patients.
$\ddagger$ Outcome homogenization is measured by pairwise consistency: the probability of an individual receiving the same outcome in 2 random allocations.
$\star$ Age homogenization is measured by Shannon entropy over the distribution of selected patients across age brackets ($\uparrow$ entropy = $\downarrow$ homogenization).
Table 3. Evaluating recovered allocations from the Rashomon set. Avg over 10 train-test splits x 25 samples of $n=1000$ patients.
5.2. Evaluating Promises of the Rashomon Set
For each Rashomon set sampling method, we compare its allocations to the full set of equal-utility allocations, revealing that the recovered allocations fall short of what the full space of equal-utility allocations offers.
In real-world healthcare allocations, Obermeyer et al. (2019) highlight the fact that selected Black patients are considerably sicker than selected White patients in terms of how many active chronic illnesses they have. This indicates discrimination based on the threshold test, which evaluates if the ratio of active chronic illnesses between selected Black and White patients is greater than 1. In addition to its unfairness, this allocation is concerning from a prioritarian point of view: there are many Black patients who were not selected who are sicker than the White patients who were selected. Across recovered allocations from the Rashomon set, we consider the least discriminatory allocation to have the smallest ratio based on the threshold test. Figure 4(a) shows that the smallest ratio recovered by all 4 methods is around 1.2, indicating that this “least discriminatory allocation” is still discriminatory. In contrast, the space of all equal-utility allocations contains many allocations with threshold test ratios at or less than 1. We specifically compare to an ideal equal-utility allocation that first selects $k^{\prime}$ qualified patients in descending order by number of chronic illnesses, then selects $k-k^{\prime}$ unqualified patients in the same fashion, alternating between Black and White patients when there are ties. As Figure 4(a) shows, this results in a least [^14] We consider this the “least discriminatory” equal-utility allocation because it adheres to the normative goal of selecting patients based on how sick they are. Technically, an equal-utility allocation with a lower ratio could be obtained by adversarially choosing the least-ill Black patients and the most-ill White patients. discriminatory equal-utility allocation with a threshold test ratio around 0.9.
Even if decision-makers use different models in the Rashomon set, outcome homogenization remains high and some qualified individuals are still systemically rejected.
We compare the pairwise consistency of individual outcomes between recovered allocations from the Rashomon set and all equal-utility allocations. Pairwise consistency represents the probability an individual would receive the same outcome under two random allocations from the set under consideration (Cooper et al., 2024). As Table 3 shows, the pairwise consistency is around 95% for recovered allocations from the Rashomon set. This is much higher than the pairwise consistency across all equal-utility allocations, which is around 75%. If decision-makers exclusively choose models from the Rashomon set, the high pairwise consistency implies that many individuals may be systemically rejected, even if they are qualified. Figure 4(b) analyzes how many qualified individuals receive multiple outcomes across recovered allocations from the Rashomon set. Depending on the method, between 10% to 30% of qualified individuals are never selected in any recovered allocation.
Consider the scenario where a fairness-oriented decision-maker uses the Rashomon set to average predictions and then selects patients with the top $k$ highest predictions from this ensemble. We show that the selected individuals under this allocation have different features from the selected individuals under a randomly-drawn equal-utility allocation. We focus on a specific feature – the age of selected patients – because of its potential to be homogenized in healthcare allocations. A stark example is the UK’s liver transplant matching algorithm, which systematically denied transplants to younger patients regardless of medical urgency (Narayanan et al., 2024). Figure 4(c) shows a similar result for the ensemble allocation from the Rashomon set: the age of selected patients is skewed older. While older patients are expected to be sicker and more qualified, Figure 4(c) also shows the age distribution of selected patients under a randomly-drawn equal-utility allocation, which chooses significantly more younger patients. Table 3 measures this phenomenon using Shannon entropy over the distribution of selected patients across age brackets. The recovered allocations from the Rashomon set have a lower entropy, indicating more homogenization in the age of selected patients.
6. Discussion
Our results show that conventional methods of using the Rashomon set do not fulfill its promises in resource-constrained allocations. Since we have established that model multiplicity is useful primarily insofar as it yields allocation multiplicity, how can we find more equal-utility allocations? We discuss three approaches below as avenues for future research.
Improving Sampling from the Rashomon Set
Allocation multiplicity should be an evaluation target for empirical methods that aim to find Rashomon models, especially given that a growing area of work involves developing and improving such methods (Rudin et al., 2024; Black et al., 2022). As Table 2 shows, all the methods that we explore find many different models in the Rashomon set. However, some methods perform better than others in terms of how many unique allocations they recover under a top $k$ mapping. We find that using different subsets of features and training data recovers more unique allocations than shuffling data order or perturbing model weights. But given that the former methods involve re-training different models from scratch, there are computational limitations to their use in practice. Even when many unique allocations are recovered, we further show that they fall short of what the full-space of equal-utility allocations promises. We hope the perspective of allocation multiplicity inspires future work to develop methods of finding diverse equally-good models that deliver on these promises.
Mitigating Structural Biases and Underspecification
Addressing structural biases and underspecification in how risk scores are generated may lead to equal-utility allocations that status-quo models fail to reflect. Even if we sample extensively from the Rashomon set, the features used to model risk may lack the richness needed to distinguish between high-risk and low-risk individuals, particularly within certain subgroups. For example, in healthcare allocation, models trained on datasets that under-represent minority populations may fail to account for systemic barriers, such as limited access to preventive care or higher exposure to environmental risks (Obermeyer et al., 2019). As a result, these individuals are often misclassified as high-risk, even when they might benefit greatly from treatment (c.f. Figure 5). This deficiency is compounded by incomplete or biased training data that limits the ability of models to learn meaningful patterns for marginalized groups. Consequently, most models in the Rashomon set may converge on overly simplistic or biased risk estimates, systematically excluding certain qualified individuals and narrowing the recovered allocation space.
Figure 5. Among equally-sick patients, Black patients have systematically lower predictions than White patients across all equally-good models.
[…]
Among equally-sick patients, Black patients have systematically lower predictions than White patients across all equally-good models.
Non-Deterministic Mappings Between Predictions and Decisions
Our case study focused on the limitations of allocations recovered by deterministically selecting individuals with the lowest risk. However, there are many stochastic procedures to map predictions to decisions (Jain et al., 2024a). Decision-makers could use a weighted lottery based on predicted risk scores, or randomize decisions for individuals near the top $k$ threshold. Table 3 shows that decision-boundary randomization outperforms all of the Rashomon set methods for each of the fairness metrics. In general, stochastic mappings may better explore the space of allocations than trying to find more models in the Rashomon set, which may not yield diverse allocations under the deterministic paradigm. While randomization involves trade-offs with expected utility, there may be little or no trade-off with realized utility, especially given the exponential number of equal-utility allocations that exist. In our case study, Table 2 shows that this trade-off is only around 1% for decision-boundary randomization, despite yielding much larger fairness benefits. We hope that the knowledge of allocation multiplicity inspires a paradigm shift in how decision-makers interpret and act on model predictions, emphasizing flexibility and equity over rigid risk-based thresholds.
Acknowledgements
We are very grateful to Prakhar Ganesh, Jamelle Watson-Daniels, and Flavio Calmon for their feedback and comments on this paper. S.J. and A.W. are generously supported by the Simons Collaboration on The Theory of Algorithmic Fairness.
Appendix
The Appendix includes the following supplementary tables and figures.
Table 4
Selected features from the Obermeyer et al. (2019) dataset used to train models in all experiments. We use the same subset of features as Watson-Daniels et al. (2023a). In particular, we do not use race as a feature.
Table 5
Performance metrics for different methods of sampling from the Rashomon set. Expanded version of Table 2, with additional results for $k/n=0.10$ and $n^{\prime}/n=0.19$, as well as standard deviation results over experiment runs.
  • The different feature subsets method has a slightly lower allocation utility compared to other methods, despite having a the lowest validation loss on average.
  • All methods have an average validation loss and deployment utility within 0.01 of each other, for all selection rates and qualification rates.
Table 6
Evaluating recovered allocations from the Rashomon set. Expanded version of Table 3, with additional results for $k/n=0.10$ and $n^{\prime}/n=0.19$, as well as standard deviation results over experiment runs.
  • Generally, the promises of the Rashomon Set are least fulfilled when scarcity is highest. In particular, the yielded benefits from the Rashomon Set are farthest from the potential benefits under all equal-utility allocations for the qualification rate of $n^{\prime}/n=0.55$.
  • For some selection rates and qualification rates, the threshold test ratio from the different feature subsets method is lower than the ratio reported for all equal-utility allocations. This is because the all equal-utility allocations ratio is computed based on the normative ideal of selecting patients based on how sick they are (in descending order by number of chronic illnesses). Technically, there exist equal-utility allocations with a lower ratio if the least-ill Black patients and most-ill White patients are selected, which the models found by the different feature subsets method may reflect.
Table 7
Additional metrics for outcome homogenization over recovered allocations from the Rashomon set. The table reports the proportion of qualified individuals that 1) are rejected in all recovered allocations, 2) receive multiple outcomes across recovered allocations, and 3) are accepted in all recovered allocations.
  • The different feature subsets and bootstrapped training data methods have the lowest rates of systemic rejection and highest rates of qualified individuals receiving multiple outcomes across the Rashomon Set.
  • When there is scarcity ($k/n<n^{\prime}/n$), the systemic rejection rate is high for all methods.
Table 8
Evaluating how stochastic mappings between predictions and decisions compare to top $k$ allocations from the Rashomon set.
  • Decision-boundary randomization refers to the partial B.F. lottery method from Jain et al. (2024a). Specifically, $\tilde{k}$ resources are randomly allocated over $\tilde{n}$ people. First, $k-\tilde{k}$ resources are deterministically given to individuals with the highest predictions. Then, an iterative weighted selection is conducted for the remaining $\tilde{k}$ resources over the next $\tilde{n}$ highest predictions. We try two parameter choices: ($\tilde{k}=0.25k,\tilde{n}=0.50k$) and ($\tilde{k}=0.50k,\tilde{n}=1.0k$).
  • Sigmoid-Logit randomization refers to a weighted lottery where the weights are determined using a sigmoid-logit transformation of predicted scores. The sigmoid-logit function (Antweiler, 2018) is defined as
    $f(x;\mu,v)=\left[1+\left(\frac{x\cdot(1-\mu)}{\mu\cdot(1-x)}\right)^{-v}\right]^{-1}$
    Figure 6 illustrates the function for different values of $\mu$ and $v$. We take $\mu=1-k/n$ and try two different values for $v$: $v=2$ and $v=5$. In a weighted lottery where the weights are based on predicted risk scores, the sigmoid-lottery function up-weights the probability that individuals in the top $k$ are selected and down-weights the probability that individuals in the bottom nkn-k are selected. Figure 6 plots the sigmoid-logit function for different values of $v$.
Figure 6. Sigmoid-logit function for varying $\mu$ and $v$ (Figure from Antweiler (2018))
Table 4. Healthcare Data Dictionary (Obermeyer et al. (2019))
Feature Description
dem_female Indicator for female gender
dem_age_band_18-24_tm1 Indicator for patient age between 18-24
dem_age_band_25-34_tm1 Indicator for patient age between 25-34
dem_age_band_35-44_tm1 Indicator for patient age between 35-44
dem_age_band_45-54_tm1 Indicator for patient age between 45-54
dem_age_band_55-64_tm1 Indicator for patient age between 55-64
dem_age_band_65-74_tm1 Indicator for patient age between 65-74
dem_age_band_75+_tm1 Indicator for patient age 75+
hypertension_elixhauser_tm1 Indicator for hypertension
cost_dialysis_tm1 Total costs for dialysis, rounded to nearest 10
cost_emergency_tm1 Total costs for emergency, rounded to nearest 10
cost_home_health_tm1 Total costs for home health, rounded to nearest 10
cost_ip_medical_tm1 Total costs for inpatient medical, rounded to nearest 10
cost_ip_surgical_tm1 Total costs for inpatient surgical, rounded to nearest 10
cost_laboratory_tm1 Total costs for laboratory, rounded to nearest 10
cost_op_primary_care_tm1 Total costs for outpatient primary care, rounded to nearest 10
cost_op_specialists_tm1 Total costs for outpatient specialists, rounded to nearest 10
cost_op_surgery_tm1 Total costs for outpatient surgery, rounded to nearest 10
cost_other_tm1 Total other costs, rounded to nearest 100
cost_pharmacy_tm1 Total costs for pharmacy, rounded to nearest 10
cost_physical_therapy_tm1 Total costs for physical therapy, rounded to nearest 10
cost_radiology_tm1 Total costs for radiology, rounded to nearest 10
gagne_sum_tm1 Total number of active illnesses
Table 5. Performance Metrics For Different Methods of Sampling From the Rashomon Set (c.f. Table 2)
Metric (Mean $\pm$ SD) $k/n$ $n^{\prime}/n$ Different Bootstrapped Shuffled Weight
Feature Subsets Training Data Data Order Perturbation
Number of Rashomon Models - 0.19 $17$ $\pm$ 3 995 $\pm$ 8 1000 $\pm$ 0 997 $\pm$ 5
- 0.32 1674 $\pm$ 3 997 $\pm$ 4 959 $\pm$ 25 999 $\pm$ 2
- 0.55 1386 $\pm$ 3 1000 $\pm$ 0 835 $\pm$ 62 1000 $\pm$ 0
Validation Loss - 0.19 0.218 $\pm$ 0.006 0.223 $\pm$ 0.006 0.222 $\pm$ 0.006 0.227 $\pm$ 0.006
- 0.32 0.279 $\pm$ 0.006 0.284 $\pm$ 0.006 0.284 $\pm$ 0.006 0.290 $\pm$ 0.006
- 0.55 0.327 $\pm$ 0.006 0.328 $\pm$ 0.007 0.329 $\pm$ 0.006 0.335 $\pm$ 0.006
Number of Recovered Allocations 0.10 0.19 570 $\pm$ 82 994 $\pm$ 8 55 $\pm$ 33 32 $\pm$ 19
0.10 0.32 652 $\pm$ 80 997 $\pm$ 4 108 $\pm$ 54 18 $\pm$ 12
0.10 0.55 691 $\pm$ 52 1000 $\pm$ 0 99 $\pm$ 32 11 $\pm$ 6
0.25 0.19 392 $\pm$ 61 994 $\pm$ 8 49 $\pm$ 24 89 $\pm$ 18
0.25 0.32 502 $\pm$ 81 997 $\pm$ 4 75 $\pm$ 46 54 $\pm$ 21
0.25 0.55 734 $\pm$ 53 1000 $\pm$ 0 184 $\pm$ 50 26 $\pm$ 11
0.50 0.19 331 $\pm$ 63 995 $\pm$ 8 75 $\pm$ 36 84 $\pm$ 17
0.50 0.32 373 $\pm$ 68 997 $\pm$ 4 62 $\pm$ 41 68 $\pm$ 11
0.50 0.55 460 $\pm$ 69 1000 $\pm$ 0 74 $\pm$ 44 31 $\pm$ 12
Allocation Utility ($k^{\prime}/k$) 0.10 0.19 0.840 $\pm$ 0.031 0.851 $\pm$ 0.032 0.852 $\pm$ 0.034 0.852 $\pm$ 0.034
0.10 0.32 0.924 $\pm$ 0.022 0.930 $\pm$ 0.024 0.931 $\pm$ 0.025 0.931 $\pm$ 0.026
0.10 0.55 0.973 $\pm$ 0.011 0.975 $\pm$ 0.013 0.975 $\pm$ 0.015 0.974 $\pm$ 0.016
0.25 0.19 0.656 $\pm$ 0.039 0.658 $\pm$ 0.039 0.659 $\pm$ 0.039 0.659 $\pm$ 0.039
0.25 0.32 0.873 $\pm$ 0.020 0.877 $\pm$ 0.020 0.877 $\pm$ 0.021 0.877 $\pm$ 0.021
0.25 0.55 0.950 $\pm$ 0.011 0.953 $\pm$ 0.012 0.952 $\pm$ 0.013 0.952 $\pm$ 0.013
0.50 0.19 0.366 $\pm$ 0.023 0.367 $\pm$ 0.023 0.367 $\pm$ 0.023 0.367 $\pm$ 0.023
0.50 0.32 0.602 $\pm$ 0.026 0.604 $\pm$ 0.026 0.605 $\pm$ 0.026 0.604 $\pm$ 0.026
0.50 0.55 0.902 $\pm$ 0.013 0.904 $\pm$ 0.013 0.905 $\pm$ 0.013 0.905 $\pm$ 0.013
Table 6. Evaluating Recovered Allocations From the Rashomon Set (c.f. Table 3)
Metric (Mean $\pm$ SD) $k/n$ $n^{\prime}/n$ All Equal-Utility Different Bootstrapped Shuffled Weight